Show that the space L^p is a normed linear space
1
Expert's answer
2020-04-04T17:26:27-0400

Let ,"1\leq p <\infin ." and "(X,\mathcal{F},\mu)" be a measurable set .Where "X" denote the under lying space , "\mathcal{F} \ the \ \sigma" -algebra of measurable set and "\mu" the measure.

Then the space "L^{p} (X,\mathcal{F},\mu)={L}^p" ,consists of all complex valued measurable function on X that satisfies


"\int_X{|f(x)|^p d\mu (x) } \ \le \infin"

The set of such function forms a vector space ,with the following natural operation

"(f+g)(x)=f(x)+g(x)"

"(\lambda f)(x)=\lambda f(x)"

For the scalar "\lambda" .

Now ,if "f\in L^p" then we define the "L^p" norm of "f" by


"\vert\vert f\mid\mid_{L^p}=(\int_X|f(x)|^pd\mu(x) )^{\frac{1}{p} }"


We also abbreviate this to "||f||_p" .

Now ,we show that "L^p" is a normed linear space under the norm "||.||_p"


Clearly ,"||f(x)||_p\geq0" ,since "|f(x)|^p\geq0" .

since "(\lambda f)x=\lambda f(x)" "\implies ||(\lambda f)x||=\lambda ||f(x)||" .

Again we known from Minkowski inequality that if "1\le p <\infin"

and "f,g\in L^p \ then \ f+g\in L^p"

and "||f+g||_p \leq ||f||_p+||g||_p" .

Hence,"L^p \ together \ with \ ||.||_p" is a semi normed vector space.

Which is denote by "{\mathcal L}^p(X,\mu)"

wayThis can be made into a normed vector space in a standard way,one simply takes the quotient space with respect to the kernel of "||.||_p" .

Since, for any measurable function f we have

"||f||_p=0 \iff f=0 \ almost \ everywhere" ,the kernel of "||.||_p"

does not depends upon p .

Let "\mathcal N =\{ f:f=0 \ almost \ everywhere \}=ker(||.||_p) , \ \forall \ 1\leq p<\infin"


In the quotient space ,two function f and g are identical if "f=g"

almost everywhere.

The resulting normed linear space is by definition,


"L^p\equiv {\mathcal L^p(X,\mu)}/{\mathcal N}"


Hence ,"L^p" is a normed linear space.